# remarkRemark

### v1

@@ 1 Introduction
The acoustic scattering problem asks for the computation of a scattered wavefield that is produced when an incoming incident wavefield strikes a scattering object. Higher-order methods for these and related waveform simulations [33] have applications in areas such virtual acoustic reconstruction [32]. When sound–soft boundary conditions are imposed on the total field $u_{tot}=u_{inc}+u$, the scattered field $u$ satisfies the following initial boundary value problem:
@@
Here, $u_{inc}$ is the incident field, $c$ is the wave speed associated with the exterior domain $\Omega$, and $[0,T]$ is some relevant time interval over which the scattered wavefield is numerically detectable. This problem is challenging to solve for a number of reasons. Standard domain discretization schemes (e.g., finite elements, finite volumes/differences) must manage the fact that the exterior domain is unbounded and impose artificial absorbing boundary layers [16] that appropriately handle outgoing waves. Traditional time-stepping methods must combat pervasive issues related to accumulating dispersive error and potentially prohibitive CFL restrictions on time step sizes. These issues are severely exacerbated when the domain includes corners that require intensive spatial refinement, and/or cavities where scattered waves decay in magnitude very slowly.
Under the condition that the incident wavefield is enveloped by a Gaussian or is otherwise approximately bandlimited, some of these complications can be avoided by using hybrid time-frequency solvers [2, 35]. In this class of solvers, the solution $u(x,t)$ is represented in terms of its Fourier transform $\hat{U}(x,\omega)$, which can be evaluated directly across a band $[W_{1},W_{2}]$ of relevant frequencies by solving the boundary value problem
@@
(5) $\hat{U}(x,\omega)=-\hat{U}_{inc}(x,\omega),\quad x\in\partial\Omega,$
where $\hat{U}_{inc}(x,\omega)$ is the Fourier transform of $u_{inc}(x,t)$ with respect to $t$. The Sommerfeld radiation condition [42] is additionally imposed to exclude incoming waves from infinity. The numerical task becomes the evaluation of the transform integral
@@
where $\hat{U}(x,\omega)$ is the solution to the Helmholtz equation at frequency $\omega$ and point $x\in\Omega$. For completeness, we note that the other half of the transform pair is
@@
Figure 1: Plots of $|u_{tot}(x,t)|$ at times $t=17.541$ (left), $t=31.311$ (center), and $t=130.00$ (right) on a “radiator” domain.
The time-frequency approach enjoys several advantages over more traditional time-domain-based methods, including the avoidance of restrictive CFL conditions, the ability to evaluate the solution at arbitrary points in time (time-skipping), and the production of solutions that are virtually free of dispersive error [2]. Moreover, issues related to discretizing the exterior domain are eliminated since solutions to the Helmholtz equation can be expressed as integrals over $\partial\Omega$. Substantial progress in the development of practical time-frequency methods has been made in [2]; related earlier work exists both in the context of Fourier transforms [35] and Laplace transforms [6, 31]. We will focus on the Fourier-based methods, which can become prohibitively expensive unless strong assumptions about the smoothness of the domain and analyticity properties of $\hat{U}(x,\omega)$ are satisfied. They struggle in domains with corners and trapping regions, which are precisely the settings where the benefits of time-frequency hybrid methods are most needed.
We introduce developments that make hybrid time-frequency methods effective in domains with these “unfriendly” features. To handle trapping, we develop a damping+correction scheme that combines contour deformation with fast transforms (the FFT and the fast sinc transform) to make evaluation of the integral in (6) tractable. This makes it possible to handle scenarios where solving the Helmholtz equation is more computationally intensive, such as in the “radiator” domain shown in Figure 1. Example domains where our method works well include domains with multiple scatterers (e.g., whisper galleries or acoustic panels), as well as domains with keyhole regions, channels, and multiple corners or cusps.
@@
Note that (P1) and (P2) are coupled: If $\tilde{x}$ is a fixed point in an activated trapping region of $\Omega$, then the decay of $|u(\tilde{x},t)|$ is slow, so we expect that large values of $t$ will be of interest. However, this is also precisely when one must worry about (P2). As shown in Figure 2, poles of the meromorphic extension111computed via the AAA algorithm [38]. of $\hat{U}(\tilde{x},\omega)$ in the complex $\omega$-plane creep closer to the real line as the trapping becomes more severe. The distance of the nearest pole to the real line corresponds to the decay rate of $|u(\tilde{x},t)|$; this phenomenon is connected to the analysis and computation of the eigenvalues and near-resonant modes of scattering operators [13, 40].
@@
where the correction terms $I_{cL}$ and $I_{cR}$ are integrals along the vertical sides of the rectangle. There is an inherent upper limit on $\delta$. If it is taken too large, the correction terms cannot be stably evaluated. However, the room afforded by $\delta$ is generous enough that when paired with a fast and accurate evaluation scheme for the highly oscillatory integral $I_{\delta}$, we are able to solve the acoustic scattering problem, even with severe trapping, over long time horizons. Combined with highly effective and well-established methods [18, 34] for solving the Helmholtz equation on challenging domains (e.g., with corners, cusps, etc.), this approach can manage complicated domains. This is illustrated with several examples in Section 5; videos of various simulations can be seen at the website listed in [46].
Our method can be applied using any (damped) Helmholtz solver that can handle complex-valued frequencies. To manage corners, one can incorporate state-of-the-art methods based on specialized series expansions [11, 24, 43], QBX-based methods [7] or more naive but easier-to-implement refinement schemes [34, Ch. 12], which is what we use in our examples. One could also consider more sophisticated strategies, such as broadband Helmholtz solvers [19]; we investigate this approach in greater depth in future work. While we emphasize 2D problems in this paper, the ideas extend naturally to the 3D setting. For readability, it should be assumed unless otherwise stated that $\Omega$ is a subset of $\mathbb{R}^{2}$.
@@ 1.1 Related work
Two papers were foundational in the development of our work. Rokhlin’s 1983 paper [41] on solving acoustic scattering problems introduces ideas fundamental to the numerical development of boundary-integral-based approaches to this problem, and also applies perturbation to the frequency parameter to alleviate trapping. Rokhlin credits the latter idea to Courant and Hilbert [14]. Several other important numerical developments for representing and computing solutions (without strong trapping) with the Fourier-based time-frequency method were introduced in Anderson, Bruno and Lyon’s 2020 paper [2], including a method for efficiently handling highly oscillatory inverse Fourier integrals via truncated sinc expansions and fast transforms. The method we apply is mathematically equivalent to this one but considers the involved ideas from the perspective of classical results on sampling schemes for the recovery of bandlimited functions.
A recently published preprint from Bruno and Santana [12] is the work closest to our own, as it tackles the same problem. Here, the authors use a polynomial + rational representation of $\hat{U}(x,\cdot)$ to formulate a singularity subtraction method [12] for handling issues with trapping. In contrast, our method uses a purely polynomial representation and applies damping to improve the generally slow convergence rate of polynomial approximation in this setting. The method in [12] is shown to be effective for a variety of challenging problems. Moreover, it has the advantage that it directly identifies near-resonances and therefore allows one to derive large-time asymptotic expansions of the scattered field. However, the approach is based on AAA approximation, and the analysis of robust convergence with respect to the number of Helmholtz solves is delicate. The ideas in [12] can likely be combined with the ideas developed here. For instance, if the resolvent-based methods in [13] for pole-finding are adapted to handle complex-valued frequencies, then the damping + correction idea can be used in concert with a polynomial + rational representation of $\hat{U}(x,\cdot)$. This would reduce both the number of poles required to accurately represent the slow-decaying components of $u(x,t)$ and the number of evaluations of $\hat{U}(x,\cdot)$ required for constructing the rational component of the approximation.
Other less-directly related works include the hybrid Fourier approach from [35] and convolutional quadrature methods based on Laplace transforms [4, 6, 5, 31]. Approaches related to retarded potential integral functions for a boundary element method in the time domain can be found in [10, 17, 49].
@@ 2 The Fourier-based hybrid time-frequency method
As seen in (5), the function $-\hat{U}_{inc}(x,\omega)$ furnishes Dirichlet boundary data. Given that $\hat{U}_{inc}$ is accessible, we consider the problem of evaluating $u(x,t)$ at all points $X\times S$, where
@@ 2.1 Transforming the incident wavefield
The class of solvers we are interested in are applicable in regimes where $u_{inc}$ decays sufficiently rapidly in both frequency space and in the time domain. This means that at any fixed point $\tilde{x}$ in space, $u_{inc}(\tilde{x},t)$ practically vanishes (e.g., $|u_{inc}(\tilde{x},t)|<\epsilon_{mach}$) outside a finite interval of support over [^2] While the scattered field zero-valued for $t<0$, this is not the case for the incident field. $t\in(-\infty,\infty)$. Additionally, the Fourier transform $\mathcal{F}(u_{inc}(\tilde{x},t))=\hat{U}_{inc}(\tilde{x},\omega)$ nearly vanishes outside a fixed band of frequencies $[W_{1},W_{2}]$. We take as a model incident field the Gaussian packet of the form
@@
A broader class of incident fields can also be considered, but the transform integrals for computing the Fourier transform of these fields are potentially highly oscillatory. A major focus of [2] is the development of a windowing method that alleviates this issue. The windowing method also results in a separation of scales in the frequency domain that can exploited in later stages of the hybrid time-frequency solver. These ideas can optionally be applied in our method, allowing for more versatility in the choice of incident waves. However, since this is not the focus of the present manuscript, we limit our consideration of incident wavefields.
@@ 2.2 The fast-decaying case
Here, $\eta_{\ell}$ and $\omega_{\ell}$ are quadrature weights and nodes, respectively. Since $t_{k}$ is never especially large and $\hat{U}(x,\cdot)$ is sufficiently nice, the size of $m$ is constrained primarily by the frequency content in the problem. The central computational challenge in computing the sum in (10) is the evaluation of $\hat{U}(x_{j},\omega_{\ell})$ over the set $\{\omega_{\ell}\}_{\ell=1}^{m}\times\{x_{j}\}_{j=1}^{M}$, which requires solving the Helmholtz equation at each fixed frequency $\omega_{\ell}$. In principle, any Helmholtz solver can be used for this purpose. A wealth of such solvers exist that can deliver high accuracy solutions even in settings where weak singularities induced by corners arise. Examples include specialized solvers that use expansions of fundamental solutions [9, 20], as well as variations of fast direct solvers that use boundary integral formulations in conjunction with specialized discretization strategies for corner geometries [11, 43] and near-singular integrals [7, 28, 1, 48]. We have implemented a solver using a boundary integral formulation, and give a brief overview of relevant details here. A more thorough review can be found in [34].
@@ 2.2.1 A boundary integral formulation
Consider a single fixed frequency $\omega_{\ell}>0$ and set $\kappa_{\ell}=\omega_{\ell}/c$. The solution to (4) at $\kappa_{\ell}$ over all $x\in\Omega$ can be expressed in terms of an integral equation that depends on an unknown density function $\phi_{\ell}$ [34, Ch. 10]. With the so-called combined field formulation, the solution is
(11) $\hat{U}(x,\omega_{\ell})=\int_{\partial\Omega}\left(d_{\kappa_{\ell}}(x,y)+iRe(\kappa_{\ell})s_{\kappa_{\ell}}(x,y)\right)\phi_{\ell}(y){\rm ds}(y),$
@@
(12) $s_{\kappa_{\ell}}(x,y)=G_{\kappa_{\ell}}(x-y),\quad d_{\kappa_{\ell}}(x,y)=\mathbf{n}(y)\cdot\nabla_{x}G_{\kappa_{\ell}}(x-y),\quad G_{\kappa_{\ell}}(x)=\frac{i}{4}H_{0}^{(1)}(\kappa_{\ell}|x|).$
Here, $H_{0}^{(1)}$ is the zeroth-order Hankel function of the first kind [39, Sec. 10.2], and $G_{\kappa_{\ell}}$ is the free-space fundamental solution to the 2D Helmholtz equation. The work required to solve (4) over $\{x_{j}\}_{j=1}^{M}\times\{\omega_{\ell}\}_{\ell=1}^{m}$ can be accomplished in two stages. First, we construct a set of density functions $\Phi=\{\phi_{1},\phi_{2},\cdots\phi_{m}\}$. Then, we use the solution formulation in (11) to evaluate each $\hat{U}(x,\omega_{\ell})$ at every $x\in X$.
The unknown function $\phi_{\ell}$ satisfies the following second-kind Fredholm equation on $L^{2}(\partial\Omega)$[34, Ch. 11]:
(13) $\frac{1}{2}\phi_{\ell}(x)+\int_{\partial\Omega}\left(d_{\kappa_{\ell}}(x,y)+i\kappa_{\ell}s_{\kappa_{\ell}}(x,y)\right)\phi_{\ell}(y){\rm ds}(y)=-\hat{U}_{inc}(x,\omega_{\ell}).$
We choose to work with the so-combined field formulation rather than those based on only the single or double layer potential because this ensures that the problem is well-posed (there are no spurious resonances on $\mathbb{R}$).
A Nyström discretization of (13) of size $n$ leads to an $n\times n$ linear system that can be solved for the vector $\tilde{\phi}_{\ell}=\begin{bmatrix}\phi_{\ell}(y_{1})&\phi_{\ell}(y_{2})&\cdots&\phi_{\ell}(y_{n})\end{bmatrix}^{T}$, which we take as a representation of $\phi_{\ell}$. Here, $n$ grows linearly with the wave number $\kappa_{\ell}$. For smooth boundaries, we apply the Zeta correction quadrature scheme from [48], which is designed to manage the singularity in the integrand of the integral operator in (11). Modifications for nonsmooth boundaries are discussed in Section 5.3. In this way, we compute and store representations for the set of density functions $\Phi=\{\phi_{1},\phi_{2},\ldots,\phi_{m}\}$.
@@
$\mathbf{x}=\begin{bmatrix}x_{1},&\cdots&x_{M}\end{bmatrix}^{T}$, by evaluating (11). If the discretization of each contour integral in (11) requires $\leq n$ quadrature points, then the naive evaluation of each integral requires a matrix vector product with an $M\times n$ matrix. The total computational cost for all $b_{\ell}(\mathbf{x})$ is $\mathcal{O}(Mnm)$. However, this cost can be reduced to only $\mathcal{O}(m(M+n))$ by instead using the fast multipole method (FMM) [23].
@@
Naively this takes an additional $\mathcal{O}(NMm)$ computations. However, since (14) has the form of a type-III nonuniform discrete Fourier transform, we can make use of the NUFFT-III [8] to do the computation in only $\mathcal{O}(M(P\log P+m+N))$ computations, where $P=W_{2}-W_{1}$. Now we turn to the setting where $|u(x,t)|$ decays slowly in time.

### v2

@@ 1 Introduction
The acoustic scattering problem asks for the computation of a scattered wavefield that is produced when an incoming incident wavefield strikes a scattering object. Higher-order methods for these and related waveform simulations [maggio1994acoustic] have applications in areas such as virtual acoustic reconstruction [lumini2023project]. When sound–soft boundary conditions are imposed on the total field $u_{tot}=u_{inc}+u$, the scattered field $u$ satisfies the following initial boundary value problem:
@@
Here, $u_{inc}$ is the incident field, $c$ is the wave speed associated with the exterior domain $\Omega$, and $[0,T]$ is some relevant time interval over which the scattered wavefield is numerically detectable. This problem is challenging to solve for a number of reasons. Standard domain discretization schemes (e.g., finite elements, finite volumes/differences) must manage the fact that the exterior domain is unbounded and impose artificial boundary conditions that mimic the behavior of the original problem (e.g. the absorbing boundary layers method of [engquist1977absorbing]). Traditional time-stepping methods must combat pervasive issues related to accumulating dispersive error and potentially prohibitive CFL restrictions on time step sizes. These issues are severely exacerbated when the domain includes corners that require intensive spatial refinement, and/or cavities where scattered waves decay in magnitude very slowly.
Under the condition that the incident wavefield is enveloped by a Gaussian or is otherwise approximately bandlimited, some of these complications can be avoided by using hybrid time-frequency solvers [anderson2020high, mecocci2000new]. In this class of solvers, the solution $u(x,t)$ is represented in terms of its Fourier transform $\hat{U}(x,\omega)$, which can be evaluated directly across a band $W_{1},W_{2}]$ of relevant frequencies by solving the boundary value problem
@@
(5) $\hat{U}(x,\omega)=-\hat{U}_{inc}(x,\omega),\quad x\in\partial\Omega,\$
where $\hat{U}_{inc}(x,\omega)$ is the Fourier transform of $u_{inc}(x,t)$ with respect to $t$. The Sommerfeld radiation condition [schot1992eighty] is additionally imposed to exclude incoming waves from infinity. The numerical task becomes the evaluation of the transform integral
@@
where $\hat{U}(x,\omega)$ is the solution to the Helmholtz equation at frequency $\omega$ and point $x\in\Omega$. Here we assume that $W_{1}>0$ is bounded away from the origin to avoid a weak singularity in $\hat{U}(x,\omega)$ as $\omega\to 0$. However, there are specialized quadrature methods that can be applied in such cases [Anderson_1984]. For completeness, we note that the other half of the transform pair is
@@
Figure 1: Plots of $|u_{tot}(x,t)|$ at times $t=17.541$ (left), $t=31.311$ (center), and $t=130.00$ (right) on a “magnetron” domain.
The time-frequency approach enjoys several advantages over more traditional time-domain-based methods, including the avoidance of restrictive CFL conditions, the ability to evaluate the solution at arbitrary points in time (time-skipping), and the production of solutions that are virtually free of dispersive error [anderson2020high]. Moreover, issues related to discretizing the exterior domain are eliminated since solutions to the Helmholtz equation can be expressed as integrals over $\partial\Omega$. Substantial progress in the development of practical time-frequency methods has been made in [anderson2020high]; related earlier work exists both in the context of Fourier transforms [mecocci2000new] and Laplace transforms [banjai2012wave, lubich1994multistep]. We will focus on the Fourier-based methods, which can become prohibitively expensive unless strong assumptions about the smoothness of the domain and analyticity properties of $\hat{U}(x,\omega)$ are satisfied. They struggle in domains with corners and trapping regions, which are precisely the settings where the benefits of time-frequency hybrid methods are most needed.
We introduce developments that make hybrid time-frequency methods effective in domains with these “unfriendly” features. To handle trapping, we develop a damping+correction scheme that combines contour deformation with fast transforms (the FFT and the fast sinc transform) to make evaluation of the integral in (6) tractable. This makes it possible to handle scenarios where solving the Helmholtz equation is more computationally intensive, such as in the “magnetron” domain shown in Figure 1. Example domains where our method works well include domains with multiple scatterers (e.g., whisper galleries or acoustic panels), as well as domains with keyhole regions, channels, and multiple corners or cusps.
@@
Note that (P1) and (P2) are coupled: If $\tilde{x}$ is a fixed point in an activated trapping region of $\Omega$, then the decay of $|u(\tilde{x},t)|$ is slow, so we expect that large values of $t$ will be of interest. However, this is also precisely when one must worry about (P2). As shown in Figure 2, poles of the meromorphic extension111computed via the AAA algorithm [nakatsukasa2018aaa]. of $\hat{U}(\tilde{x},\omega)$ in the complex $\omega$-plane creep closer to the real line as the trapping becomes more severe. The distance of the nearest pole to the real line corresponds to the decay rate of $|u(\tilde{x},t)|$; this phenomenon is connected to the analysis and computation of the eigenvalues and near-resonant modes of scattering operators [bruno2024evaluation, patel2023wigner].
@@
where the correction terms $I_{cL}$ and $I_{cR}$ are integrals along the vertical sides of the rectangle. There is an inherent upper limit on $\delta$. If it is taken too large, the correction terms cannot be stably evaluated. However, the room afforded by $\delta$ is generous enough that when paired with a fast and accurate evaluation scheme for the highly oscillatory integral $I_{\delta}$, we are able to solve the acoustic scattering problem, even with severe trapping, over long time horizons. Combined with highly effective and well-established methods [gillman2012direct, martinsson2019fast] for solving the Helmholtz equation on challenging domains (e.g., with corners, cusps, etc.), this approach can manage complicated domains. This is illustrated with several examples in Section 5; videos of various simulations can be seen at the website listed in [HOMEyoutube].
Our method can be applied using any (damped) Helmholtz solver that can handle complex-valued frequencies. To manage corners, one can incorporate state-of-the-art methods based on specialized series expansions [bruno2009high, hoskins2019numerical, serkh2019solution], QBX-based methods [barnett2014evaluation] or more naive but easier-to-implement refinement schemes [martinsson2019fast, Ch. 12], which is what we use in our examples. One could also consider more sophisticated strategies, such as broadband Helmholtz solvers [gopal2022broadband]; we investigate this approach in greater depth in future work. While we emphasize 2D problems in this paper, the ideas extend naturally to the 3D setting (see Figure 10). For readability, it should be assumed unless otherwise stated that $\Omega$ is a subset of $\mathbb{R}^{2}$.
@@ 1.1 Related work
Two papers were foundational in the development of our work. Rokhlin’s 1983 paper [rokhlin1983solution] on solving acoustic scattering problems introduces ideas fundamental to the numerical development of boundary-integral-based approaches to this problem, and also applies perturbation to the frequency parameter to alleviate trapping. Rokhlin credits the latter idea to Courant and Hilbert [courant1962methods]. Several other important numerical developments for representing and computing solutions (without strong trapping) with the Fourier-based time-frequency method were introduced in Anderson, Bruno and Lyon’s 2020 paper [anderson2020high], including a method for efficiently handling highly oscillatory inverse Fourier integrals via truncated sinc expansions and fast transforms. The method we apply is mathematically equivalent to this one but considers the involved ideas from the perspective of classical results on sampling schemes for the recovery of bandlimited functions.
A recently published preprint from Bruno and Santana [bruno2025efficient] is the work closest to our own, as it tackles the same problem. Here, the authors use a polynomial + rational representation of $\hat{U}(x,\cdot)$ to formulate a singularity subtraction method [bruno2025efficient] for handling issues with trapping. In contrast, our method uses a purely polynomial representation and applies damping to improve the generally slow convergence rate of polynomial approximation in this setting. The method in [bruno2025efficient] is shown to be effective for a variety of challenging problems. Moreover, it has the advantage that it directly identifies near-resonances and therefore allows one to derive large-time asymptotic expansions of the scattered field. However, the approach is based on AAA approximation, and the analysis of robust convergence with respect to the number of Helmholtz solves is delicate. The ideas in [bruno2025efficient] can likely be combined with the ideas developed here. For instance, if the resolvent-based methods in [bruno2024evaluation] for pole-finding are adapted to handle complex-valued frequencies, then the damping + correction idea can be used in concert with a polynomial + rational representation of $\hat{U}(x,\cdot)$. This would reduce both the number of poles required to accurately represent the slow-decaying components of $u(x,t)$ and the number of evaluations of $\hat{U}(x,\cdot)$ required for constructing the rational component of the approximation.
Other less-directly related works include the hybrid Fourier approach from [mecocci2000new] and convolutional quadrature methods based on Laplace transforms [banjai2010multistep, banjai2012wave, banjai2014fast, lubich1994multistep]. Approaches related to retarded potential integral functions for a boundary element method in the time domain can be found in [barnett2020high, et1986formulation, yilmaz2004time].
@@ 2.1 Transforming the incident wavefield
As seen in (5), the incoming field $\hat{U}_{inc}(x,\omega)$ furnishes Dirichlet boundary data. The class of solvers we are interested in are applicable in regimes where $u_{inc}$ decays sufficiently rapidly in both frequency space and in the time domain. This means that at any fixed point $\tilde{x}$ in space, $u_{inc}(\tilde{x},t)$ practically vanishes (e.g., $|u_{inc}(\tilde{x},t)|<\epsilon_{mach}$) outside a finite interval of support over [^2] While the scattered field is zero-valued for $t<0$, this is not the case for the incident field. $t\in(-\infty,\infty)$. Additionally, the Fourier transform $\hat{U}_{inc}(\tilde{x},\omega)=\mathcal{F}[u_{inc}(\tilde{x},\cdot)](\omega)$, nearly vanishes outside a fixed band of frequencies $W_{1},W_{2}]$. We take as a model incident field the Gaussian packet of the form
@@
A broader class of incident fields can also be considered, but the transform integrals for computing the Fourier transform of these fields are potentially highly oscillatory. A major focus of [anderson2020high] is the development of a windowing method that alleviates this issue. The windowing method also results in a separation of scales in the frequency domain that can exploited in later stages of the hybrid time-frequency solver. These ideas can optionally be applied in our method, allowing for more versatility in the choice of incident waves. However, since this is not the focus of the present manuscript, we limit our consideration of incident wavefields. Given that $\hat{U}_{inc}$ is accessible, we consider the problem of evaluating $u(x,t)$ at all points $X\times S$, where
@@ 2.2 The fast-decaying case
Here, $\eta_{\ell}$ and $\omega_{\ell}$ are quadrature weights and nodes, respectively. Since $t_{k}$ is never especially large and $\hat{U}(x,\cdot)$ is sufficiently nice, the size of $m$ is constrained primarily by the frequency content in the problem. The central computational challenge in computing the sum in (10) is the evaluation of $\hat{U}(x_{j},\omega_{\ell})$ over the set $\{\omega_{\ell}\}_{\ell=1}^{m}\times\{x_{j}\}_{j=1}^{M}$, which requires solving the Helmholtz equation at each fixed frequency $\omega_{\ell}$. In principle, any Helmholtz solver can be used for this purpose. A wealth of such solvers exist that can deliver high accuracy solutions even in settings where weak singularities induced by corners arise. Examples include specialized solvers that use expansions of fundamental solutions [barnett2010exponentially, gopal2019], as well as variations of fast direct solvers that use boundary integral formulations in conjunction with specialized discretization strategies for corner geometries [bruno2009high, serkh2019solution] and near-singular integrals [af2018adaptive, barnett2014evaluation, krantz2024error, wu2021zeta]. We have implemented a solver using a boundary integral formulation, and give a brief overview of relevant details here. A more thorough review can be found in [martinsson2019fast].
@@ 2.2.1 A boundary integral formulation
Consider a single fixed frequency $\omega_{\ell}>0$ and set $\kappa_{\ell}=\omega_{\ell}/c$. The solution to (4) at $\kappa_{\ell}$ over all $x\in\Omega$ can be expressed in terms of an integral equation that depends on an unknown density function $\phi_{\ell}$ [martinsson2019fast, Ch. 10]. With the so-called combined field formulation, the solution is
(11) $\hat{U}(x,\omega_{\ell})=\int_{\partial\Omega}\left(d_{\kappa_{\ell}}(x,y)-iRe(\kappa_{\ell})s_{\kappa_{\ell}}(x,y)\right)\phi_{\ell}(y){\rm ds}(y),$
@@
(12) $s_{\kappa_{\ell}}(x,y)=G_{\kappa_{\ell}}(x-y),\quad d_{\kappa_{\ell}}(x,y)=n(y)\cdot\nabla_{x}G_{\kappa_{\ell}}(x-y),\quad G_{\kappa_{\ell}}(x)=\frac{i}{4}H_{0}^{(1)}(\kappa_{\ell}|x|).$
Here, $H_{0}^{(1)}$ is the zeroth-order Hankel function of the first kind [olver2010nist, Sec. 10.2], $n(y)$ is the outward normal unit vector to $\partial\Omega$, and $G_{\kappa_{\ell}}$ is the free-space fundamental solution to the 2D Helmholtz equation. The work required to solve (4) over $\{x_{j}\}_{j=1}^{M}\times\{\omega_{\ell}\}_{\ell=1}^{m}$ can be accomplished in two stages. First, we construct a set of density functions $\Phi=\{\phi_{1},\phi_{2},\cdots\phi_{m}\}$. Then, we use the solution formulation in (11) to evaluate each $\hat{U}(x,\omega_{\ell})$ at every $x\in X$.
The unknown function $\phi_{\ell}$ satisfies the following second-kind Fredholm equation on $L^{2}(\partial\Omega)$[martinsson2019fast, Ch. 11]:
(13) $\frac{1}{2}\phi_{\ell}(x)+\int_{\partial\Omega}\left(d_{\kappa_{\ell}}(x,y)-i\kappa_{\ell}s_{\kappa_{\ell}}(x,y)\right)\phi_{\ell}(y){\rm ds}(y)=-\hat{U}_{inc}(x,\omega_{\ell}).$
We choose to work with the “combined field formulation” rather than those based on only the single or double layer potential because this ensures that the problem is well-posed (there are no spurious resonances).
A Nyström discretization of (13) of size $n$ leads to an $n\times n$ linear system that can be solved for the vector $\tilde{\phi}_{\ell}=\begin{bmatrix}\phi_{\ell}(y_{1})&\phi_{\ell}(y_{2})&\cdots&\phi_{\ell}(y_{n})\end{bmatrix}^{T}$, which we take as a representation of $\phi_{\ell}$. Here, $n$ grows linearly with the wave number $\kappa_{\ell}$. For smooth boundaries, we apply the Zeta correction quadrature scheme from [wu2021zeta], which is designed to manage the singularity in the integrand of the integral operator in (11). Modifications for nonsmooth boundaries are discussed in Section 5.3. In this way, we compute and store representations for the set of density functions.
@@
$\mathbf{x}=\begin{bmatrix}x_{1},&\cdots&x_{M}\end{bmatrix}^{T}$, by evaluating (11). If the discretization of each contour integral in (11) requires $\leq n$ quadrature points, then the naive evaluation of each integral requires a matrix vector product with an $M\times n$ matrix. The total computational cost for all $b_{\ell}(\mathbf{x})$ is $\mathcal{O}(Mnm)$. However, this cost can be reduced to only $\mathcal{O}(m(M+n))$ by instead using the fast multipole method (FMM) [greengard1987fast].
@@
Naively this takes an additional $\mathcal{O}(NMm)$ computations. However, since (14) has the form of a type-III nonuniform discrete Fourier transform, we can make use of the NUFFT-III [barnett2018non] to do the computation in only $\mathcal{O}(M(P\log P+m+N))$ computations, where $P=W_{2}-W_{1}$. Now we turn to the setting where $|u(x,t)|$ decays slowly in time.